### v2

@@ 4.2.6 Model-Agnostic Frameworks
Table 7: Taxonomy and Methodologies of Traditional Deep Learning Architectures for Time Series Forecasting
@@
    Koopa MLP · Koopa Block with Koopman Predictor(KP) CD 2023
@@
    WaveForM Hybrid (CNN+GCN) · Discrete Wavelet Transform(DWT) Module · Graph-Enhanced Prediction Module CD 2023
    BiTGraph · Multi-Scale Instance PartialTCN(MSIPT) Module · Biased GCN Module CD 2024
    TSLANet Hybrid (TF+CNN) · Adaptive Spectral Block · Interactive Convolutional Block CI 2024
    DERITS Hybrid (CNN+MLP) · Frequency Derivative Transformation(FDT) · Order-adaptive Fourier Convolution Network(OFCN) CD 2024
@@ 4.3 Emergence of Foundation Models
In recent years, foundation models have demonstrated remarkable performance and strong zero-shot capabilities across various tasks and domains, leading to increased focus [172, 108, 95, 2, 145]. While significant progress has been made in domains like vision and language, developing foundation models for time series data has faced several challenges. Firstly, time series data exhibit distinct characteristics depending on the dataset. For instance, electrocardiogram (ECG), weather, and sensor data from industrial processes have unique properties in terms of variables, frequency, periodicity, and noise, often requiring domain-specific knowledge for effective modeling. Additionally, unlike the vision and language domains, where large-scale pre-training corpora can be relatively easily constructed from publicly available sources like the web, collecting time series data is more difficult due to high acquisition costs and security concerns. Despite these obstacles, research on time series foundation models is essential for improving model scalability and generality. This necessity has led to active exploration in several key directions, particularly in the field of time series forecasting.
@@ 4.3.1 Sequence Modeling with LLMs
Both GPT4TS and PromptCast rely on fine-tuning to achieve their results. In contrast, without fine-tuning, LLMTime [64] has demonstrated impressive zero-shot forecasting capabilities. The main idea of LLMTime is encoding time series data as a string of numerical digits, which emphasizes the importance of preprocessing and its inherent dependencies. To align the time series modality with the language modality, Time-LLM [85] introduces the concept of patch reprogramming. This concept aims to mitigate the challenges of a large reprogramming space and attempts to connect time series local characteristics with language semantics, such as “short up”. While the patch reprogramming offers more flexibility by reducing the reliance on large-scale time series corpora, it also presents the challenge of adapting time series data to fit the characteristics of large language models.
@@ 4.3.2 Pre-training
Another approach that focuses on building large-scale time series corpora to pre-train time series foundation models from scratch has emerged. An example is Lag-LLaMA [148], which uses a decoder-only Transformer to generate forecasting results, enabling probabilistic forecasting. Lag-LLaMA also consolidates 27 publicly available forecasting datasets from various domains to create a comprehensive pre-training corpus. On the other hand, TimesFM [36] extends beyond publicly available forecasting datasets by incorporating additional pre-training corpora based on Google Trends and Wiki Pageview statistics. It also utilizes synthetic datasets generated from piece-wise linear trends, autoregressive processes, and sine and cosine functions to capture universal characteristics. The entire pre-training corpus spans hundreds of billions of time steps. While most of the time series foundation models rely on temporal embedding, CHRONOS [7] takes a different idea by learning a patch-based tokenizer, similar to traditional language models, to capture the intrinsic “language” of time series data.
Traditional foundation models often overlook the relationships between variables in multivariate time series forecasting, typically extending to multivariate forecasting by independently combining univariate forecasts based on channel independence. Uni2TS [190] addresses this limitation by explicitly considering the expansion to arbitrary multivariate time series using variate IDs. Additionally, it leverages a large-scale time series dataset called LOTSA, which accounts for multi-domain and multi-frequency characteristics.
@@ 4.4.2 Time Series Feature Extraction
Decomposition techniques involve breaking down time series data into components such as trend, seasonality, and irregularity, analyzing the unique patterns of each component. Diffusion-TS [208] points out that traditional methods cannot properly learn each component because the forward process causes the components to collapse. Therefore, it models the decomposed time series data individually during the diffusion process, learning each component, such as trend, seasonality, and residuals, independently and then recombining them to restore the entire time series data.
@@
Fourier analysis, a type of decomposition technique, converts time series data into the frequency domain to analyze periodic components. This method helps identify periodic patterns and remove noise. [33] explore the idea that the representation of time series data in the frequency domain can provide useful inductive biases for score-based diffusion models. This paper demonstrates that the components of time series data are more clearly represented in the frequency domain, and diffusion models in frequency domain outperform those in the time domain.
@@ 4.4.3 Additional Approaches
Score-Based Generative Modeling through Stochastic Differential Equations (SDEs) [163] serves as continuous energy-based generative models, handling data in a continuous time domain, thus reflecting natural and precise changes and noise in the real world. The previously discussed methods are based on DDPM [70], which add and remove noise incrementally using fixed-time steps. However, the DDPM approach faces challenges such as specific functional constraints and sensitivity to hyperparameters. Some research applying SDEs overcomes these limitations and explores diverse approaches. ScoreGrad [200] is the first work to apply SDE, overcoming the constraints of DDPM-based models and offering a more flexible and powerful generative model. D3M [201] addresses the limitations of traditional SDEs, such as the complexity of determining drift and diffusion coefficients and slow generation speed, by utilizing a decomposable noise-removal diffusion model based on explicit solutions. This method reduces computational complexity through clear SDE formulations and separates the signal decay and noise injection processes in model design. As a result, it enhances model efficiency and accuracy while accelerating generation speed. [33] improve denoising score matching in the frequency domain by using mirrored Brownian motions instead of standard Brownian motion, emphasizing the symmetry between components when applying SDEs.
@@
Some works use guidance instead of explicitly feeding conditions into the denoising network for forecasting. Diffusion-TS [208] employs classifier guidance, using a separate classifier to guide the sampling process through the gradients of classifier. This method maintains the basic unconditional diffusion model while performing conditional generation tasks through various classifiers. It generates samples that are better aligned with specific conditions, resulting in higher-quality outputs. However, classifier guidance requires a classifier for each time step, which necessitates training a new classifier instead of using pre-trained ones. LDT [47] uses classifier-free guidance, learning both conditional and unconditional models within a single model to perform conditional sampling. This work eliminates the need for an additional classifier and implicitly obtains a classifier, making implementation simpler and more efficient. TSDiff [91] proposes observation self-guidance, allowing the use of an unconditional diffusion model for conditional generation without separate networks or modifications to the training process.
@@ 4.5 Debut of the Mamba
4.5.1 History of the SSM(State Space Model)
@@
SSM is a mathematical framework used to model the dynamic state of a system that changes over time, compressing only the essential information for effective sequential modeling. SSM describes the relationship between the internal state of the system and external observations and is used in various fields such as control theory, signal processing, and time series analysis. It comprises a ‘State equation’ that explains how the internal state changes over time and an ‘Observation equation’ that explains how the internal state is transformed into external observations. Although it is a continuous model that performs linear transformations on the current hidden state and input, it can handle discrete sequences like time series through discretization. It closely resembles RNNs in that it combines observation data and hidden state data.
@@
Early applications of SSMs, such as the S4 [67] model, utilized diagonalization of the transition matrix to effectively model long-term dependencies in long and complex sequence data. These models were able to achieve high performance when combined with deep learning architectures like Transformers, RNNs, and CNNs. Subsequently, based on S4, advanced blocks like H3 [49] were developed, which hierarchically structured convolution, gating, and MLP to provide more efficient and powerful sequence modeling.
@@ 4.5.2 Introduction of the Mamba
Mamba addresses these limitations of SSMs by introducing a advanced deep structured SSM model with selective SSM. It is designed so that the parameters of the SSM dynamically change depending on the input, allowing for selective memory or disregard of specific parts of the sequence, enabling efficient data processing. However, this approach, while enhancing the system’s flexibility to better learn complex patterns, sacrifices the parallel processing advantages of SSMs. To compensate, traditional techniques such as kernel fusion, parallel scan, and recomputation are applied to efficiently implement selective SSMs on GPUs. Mamba adopts a simplified architecture centered around selective SSM, replacing the first multiplicative gate in the traditional H3 block with an activation function and incorporating an SSM into the gated MLP block. Due to the absence of attention mechanisms or separate MLP blocks, computational complexity is reduced, resulting in efficient training, fast inference, and high scalability and versatility.
@@ 5.2 Alleviation of Distribution Shift
This survey focuses on one prominent approach that introduces normalization and denormalization modules within a normalization-TSF model-denormalization framework. Here, normalization is applied to the look-back window before feeding it into the TSF model, aiming to remove input statistics. As a result, the TSF model processes inputs from a less time-variant distribution. Denormalization is then applied to the forecasting window obtained from the TSF model to restore the original statistics. The denormalized forecasting window is ultimately used as the final forecast. However, calculating the appropriate statistics for normalization and denormalization in non-stationary scenarios is a non-trivial task, and numerous studies have been proposed within this framework to address this challenge.
@@ 5.3 Enhancing Causality
Moreover, causal analysis enhances the interpretability and practical application of forecasting models. By explaining how different variables influence the outcome, causal models provide valuable insights into the mechanisms driving the forecast, which is crucial for decision-making. Businesses and policymakers can use these insights to predict the impact of specific actions, such as policy changes or marketing strategies, and make informed decisions based on the likely effects. This ability to simulate interventions and conduct counterfactual senarios makes causal analysis an indispensable tool for accurate and actionable time series forecasting.
@@ 5.4.1 Decomposition
Autoformer [191] goes beyond using decomposition techniques merely as preprocessing for forecasting tasks by progressively decomposing time series data throughout the prediction process within the model itself. Similarly, CrossWaveNet [80]employs a dual-channel network to perform gradual deep cross-decomposition, enabling it to capture complex temporal patterns effectively. Likewise, models such as FEDformer [220], LTSF-Linear [209], and PDMLP [169] utilize a moving average kernel to individually model each component of the time series data, which are then recombined to make effective predictions. However, the Moving Average method lacks robustness because it is not learnable by the model. Additionally, since it assigns equal weights to each data point within the sliding window, it has limited ability to identify specific patterns. To address these limitations, Leddam [207] use learnable 1D convolutional kernels, which can better capture nonlinear structures and dynamic variations. Meanwhile, diffusion-based models face challenges as components like trend and seasonality can easily collapse during the diffusion process. Models like Diffusion-TS [208] overcome this by applying the diffusion process after decomposition, thereby preserving the characteristics of each component more effectively. Through this approach, the traditional limitations of time series decomposition are overcome, allowing for a more effective separation of trend and seasonal components.

### v3

@@ 4.2.5 Model-Agnostic Frameworks
Table 7: Taxonomy and Methodologies of Fundamental Deep Learning Architectures for Time Series Forecasting
@@
    N-BEATS MLP · Neural Basis Expansion (Polynomial Basis & Fourier Basis) CI 2020
    N-HiTS · Neural Hierarchical Interpolation · Multi-Rate Data Sampling CI 2023
    Koopa · Koopa Block with Koopman Predictor(KP) CD 2023
@@ 4.3 Emergence of Foundation Models
In recent years, foundation models have demonstrated remarkable performance and strong zero-shot capabilities across various tasks and domains, leading to increased focus [243, 148, 131, 2, 69, 201]. While significant progress has been made in domains like vision and language, developing foundation models for time series data has faced several challenges. Firstly, time series data exhibit distinct characteristics depending on the dataset. For instance, electrocardiogram (ECG), weather, and sensor data from industrial processes have unique properties in terms of variables, frequency, periodicity, and noise, often requiring domain-specific knowledge for effective modeling. Additionally, unlike the vision and language domains, where large-scale pre-training corpora can be relatively easily constructed from publicly available sources like the web, collecting time series data is more difficult due to high acquisition costs and security concerns. Despite these obstacles, research on time series foundation models is essential for improving model scalability and generality. This necessity has led to active exploration in several key directions, particularly in the field of time series forecasting.
@@ 4.3.1 Sequence Modeling with LLMs
Both GPT4TS and PromptCast rely on fine-tuning to achieve their results. In contrast, without fine-tuning, LLMTime [87] has demonstrated impressive zero-shot forecasting capabilities. The main idea of LLMTime is encoding time series data as a string of numerical digits, which emphasizes the importance of preprocessing and its inherent dependencies. To align the time series modality with the language modality, Time-LLM [114] introduces the concept of patch reprogramming. This concept aims to mitigate the challenges of a large reprogramming space and attempts to connect time series local characteristics with language semantics, such as “short up”. While patch reprogramming offers more flexibility by reducing the reliance on large-scale time series corpora, it also presents the challenge of adapting time series data to fit the characteristics of large language models.
@@ 4.3.2 Pre-training
Another approach that focuses on building large-scale time series corpora to pre-train time series foundation models from scratch has emerged. An example is Lag-LLaMA [204], which uses a decoder-only Transformer to generate forecasting results, enabling probabilistic forecasting. Lag-LLaMA also consolidates 27 publicly available forecasting datasets from various domains to create a comprehensive pre-training corpus. On the other hand, TimesFM [48] extends beyond publicly available forecasting datasets by incorporating additional pre-training corpora based on Google Trends and Wiki Pageview statistics. It also utilizes synthetic datasets generated from piece-wise linear trends, autoregressive processes, and sine and cosine functions to capture universal characteristics. The entire pre-training corpus spans hundreds of billions of time steps. While most of the time series foundation models rely on temporal embedding, CHRONOS [7] takes a different idea by learning a patch-based tokenizer, similar to conventional language models, to capture the intrinsic “language” of time series data.
Conventional foundation models often overlook the relationships between variables in multivariate time series forecasting, typically extending to multivariate forecasting by independently combining univariate forecasts based on channel independence. Uni2TS [266] addresses this limitation by explicitly considering the expansion to arbitrary multivariate time series using variate IDs. Additionally, it leverages a large-scale time series dataset called LOTSA, which accounts for multi-domain and multi-frequency characteristics.
@@ 4.4 Advance of Diffusion Models
Diffusion models, theoretically proposed by [227], are inspired by non-equilibrium statistical physics. They learn by progressively adding noise to the data and then gradually reversing the process to recover the original data. However, this initial research lacked concrete methods for training, making practical implementation challenging. This issue was addressed by Denoising Diffusion Probabilistic Models (DDPM) [97] and Noise Conditional Score Networks (NCSN) [229]. DDPM models the forward and reverse processes explicitly based on a probabilistic approach, enabling efficient learning. NCSN, on the other hand, employs score matching techniques to refine the noise removal process, resulting in the generation of higher-quality samples.
@@
Diffusion models have a significant advantage in modeling uncertainty through the forward and reverse processes. By providing multiple possible prediction outcomes instead of a single prediction, it can reflect the uncertainty of the real world, making it highly beneficial for TSF. Therefore, it can offer prediction confidence intervals or enable probabilistic forecasting, thereby contributing to decision-making.
@@ 4.4.2 Time Series Feature Extraction
Decomposition techniques involve breaking down time series data into components such as trend, seasonality, and irregularity, analyzing the unique patterns of each component. Diffusion-TS [293] points out that conventional methods cannot properly learn each component because the forward process causes the components to collapse. Therefore, it models the decomposed time series data individually during the diffusion process, learning each component, such as trend, seasonality, and residuals, independently and then recombining them to restore the entire time series data.
@@
Fourier analysis, a type of decomposition technique, converts time series data into the frequency domain to analyze periodic components. This method helps identify periodic patterns and remove noise. [44] explore the idea that the representation of time series data in the frequency domain can provide useful inductive biases for score-based diffusion models. This paper demonstrates that the components of time series data are more clearly represented in the frequency domain, and diffusion models in the frequency domain outperform those in the time domain.
@@ 4.4.3 Additional Approaches
Score-Based Generative Modeling through Stochastic Differential Equations (SDEs) [230] serves as continuous energy-based generative models, handling data in a continuous time domain, thus reflecting natural and precise changes and noise in the real world. The previously discussed methods are based on DDPM [97], which add and remove noise incrementally using fixed-time steps. However, the DDPM approach faces challenges such as specific functional constraints and sensitivity to hyperparameters. Some research applying SDEs overcomes these limitations and explores diverse approaches. ScoreGrad [282] is the first work to apply SDE, overcoming the constraints of DDPM-based models and offering a more flexible and powerful generative model. D3M [283] addresses the limitations of conventional SDEs, such as the complexity of determining drift and diffusion coefficients and slow generation speed, by utilizing a decomposable noise-removal diffusion model based on explicit solutions. This method reduces computational complexity through clear SDE formulations and separates the signal decay and noise injection processes in model design. As a result, it enhances model efficiency and accuracy while accelerating generation speed. [44] improve denoising score matching in the frequency domain by using mirrored Brownian motions instead of standard Brownian motion, emphasizing the symmetry between components when applying SDEs.
@@
Some works use guidance instead of explicitly feeding conditions into the denoising network for forecasting. Diffusion-TS [293] employs classifier guidance, using a separate classifier to guide the sampling process through the gradients of the classifier. This method maintains the basic unconditional diffusion model while performing conditional generation tasks through various classifiers. It generates samples that are better aligned with specific conditions, resulting in higher-quality outputs. However, classifier guidance requires a classifier for each time step, which necessitates training a new classifier instead of using pre-trained ones. LDT [64] uses classifier-free guidance, learning both conditional and unconditional models within a single model to perform conditional sampling. This work eliminates the need for an additional classifier and implicitly obtains a classifier, making implementation simpler and more efficient. TSDiff [127] proposes observation self-guidance, allowing the use of an unconditional diffusion model for conditional generation without separate networks or modifications to the training process.
@@ 4.5 Debut of the Mamba
4.5.1 History of the State Space Models (SSMs)
@@
SSM is a mathematical framework used to model the dynamic state of a system that changes over time, compressing only the essential information for effective sequential modeling. SSM describes the relationship between the internal state of the system and external observations and is used in various fields such as control theory, signal processing, and time series analysis. It comprises a ‘State equation’ that explains how the internal state changes over time and an ‘Observation equation’ that explains how the internal state is transformed into external observations. Although it is a continuous model that performs linear transformations on the current hidden state and input, it can handle discrete sequences like time series through discretization (Fig. 13). It closely resembles RNNs in that it combines observation data and hidden state data.
@@
Early applications of SSMs, such as the S4 [89] model, utilized diagonalization of the transition matrix to effectively model long-term dependencies in long and complex sequence data. These models were able to achieve high performance when combined with deep learning architectures like Transformers, RNNs, and CNNs. Subsequently, based on S4, advanced blocks like H3 [67] were developed, which hierarchically structured convolution, gating, and MLP to provide more efficient and powerful sequence modeling (Fig. 14).
@@ 4.5.2 Introduction of the Mamba
Mamba addresses these limitations of SSMs by introducing an advanced deep structured SSM model with selective SSM. It is designed so that the parameters of the SSM dynamically change depending on the input, allowing for selective memory or disregard of specific parts of the sequence, enabling efficient data processing. However, this approach, while enhancing the system’s flexibility to better learn complex patterns, sacrifices the parallel processing advantages of SSMs. To compensate, traditional techniques such as kernel fusion, parallel scan, and recomputation are applied to efficiently implement selective SSMs on GPUs. Mamba adopts a simplified architecture centered around selective SSM, replacing the first multiplicative gate in the traditional H3 block with an activation function and incorporating an SSM into the gated MLP block. Due to the absence of attention mechanisms or separate MLP blocks, computational complexity is reduced, resulting in efficient training, fast inference, and high scalability and versatility.
@@ 5.2 Alleviation of Distribution Shift
Representative approaches for addressing distribution shifts include Domain Adaptation, Transfer Learning, and Robustness Techniques. Domain Adaptation aims to reduce the distribution differences between the source and target domains to improve generalization performance across domains. This approach is divided into supervised domain adaptation, which uses some labeled data from the target domain, and unsupervised domain adaptation which aligns feature distributions using unlabeled datasets. Unsupervised domain adaptation began with the introduction of Maximum Mean Discrepancy (MMD) [254] -based methods and advanced with the emergence of Generative Adversarial Networks (GANs) [213], which led to the development of adversarial domain adaptation techniques. Transfer Learning [182] refers to applying a model trained in one domain or task to a new dataset. This includes Feature Transfer, which reuses or retrains specific layers of a pre-trained network for a new task, and Fine-Tuning which adjusts the entire pre-trained model to fit the new dataset. Transfer learning originated from pre-trained embeddings in NLP in the late 1990s, expanded through CNNs, and eventually evolved with LLMs. Robustness Techniques enhance a model’s ability to maintain performance despite data uncertainties, such as noise, outliers, and data scarcity. Methods such as Dropout [233] and Data Augmentation [265] help models handle noise, while robust loss functions like the Huber loss [108] reduce sensitivity to outliers. These techniques have been widely adopted in time series forecasting models since 2020.
This survey focuses on one prominent approach that introduces normalization and denormalization modules within a normalization-TSF model-denormalization framework. This framework has evolved as a specialized approach for mitigating non-stationarity within the same domain of time series data. Here, normalization is applied to the look-back window before feeding it into the TSF model, aiming to remove input statistics. As a result, the TSF model processes inputs from a less time-variant distribution. Denormalization is then applied to the forecasting window obtained from the TSF model to restore the original statistics. The denormalized forecasting window is ultimately used as the final forecast. However, calculating the appropriate statistics for normalization and denormalization in non-stationary scenarios is a non-trivial task, and numerous studies have been proposed within this framework to address this challenge.
@@ 5.3 Enhancing Causality
Moreover, causal analysis enhances the interpretability and practical application of forecasting models. By explaining how different variables influence the outcome, causal models uncover the underlying mechanisms driving the forecast, which is crucial for decision-making. Businesses and policymakers can use these insights to predict the impact of specific actions, such as policy changes or marketing strategies, and make informed decisions based on the likely effects. This ability to simulate interventions and conduct counterfactual scenarios makes causal analysis an indispensable tool for accurate and actionable time series forecasting.
@@
In TSF, various causal inference methods are employed to identify cause-and-effect relationships within the data. The Granger Causality Test [85] examines whether the historical information of one variable helps predict the future of another variable. This method detects the direction of causal influence based on regression analysis. However, it has limitations in fully excluding indirect correlations caused by the presence of a third variable. Structural Causal Models (SCM) [192] utilize causal graphs and structural equations to model the relationships between variables. These models enable intervention simulations and counterfactual analyses, providing a visual representation of causal relationships to aid interpretation and integrate interactions across multiple variables. Do-Calculus [190] is an intervention analysis technique for quantitatively analyzing the effects of interventions by computing the impact that changes in specific variables have on others, thereby supporting causal predictions. Propensity Score Matching (PSM) [207] is a method for performing causal inference by matching groups with similar characteristics. This approach minimizes the influence of confounders and more accurately evaluates the effects of interventions. Directed Acyclic Graphs (DAGs) [191] are directional, non-cyclic graphs that visually represent the relationships between causes and effects. DAGs help clearly identify causal structures and understand complex interactions among variables. Various attempts are being made to enhance the causality of TSF by utilizing these diverse causal inference methods [210, 215].
@@ 5.4.1 Decomposition
Autoformer [269] goes beyond using decomposition techniques merely as preprocessing for forecasting tasks by progressively decomposing time series data throughout the prediction process within the model itself. Similarly, CrossWaveNet [107] employs a dual-channel network to perform gradual deep cross-decomposition, enabling it to capture complex temporal patterns effectively. Likewise, models such as FEDformer [311], LTSF-Linear [294], and PDMLP [239] utilize a moving average kernel to individually model each component of the time series data, which are then recombined to make effective predictions. However, the Moving Average method lacks robustness because it is not learnable by the model. Additionally, since it assigns equal weights to each data point within the sliding window, it has limited ability to identify specific patterns. To address these limitations, Leddam [292] use learnable 1D convolutional kernels, which can better capture nonlinear structures and dynamic variations. Meanwhile, diffusion-based models face challenges as components like trend and seasonality can easily collapse during the diffusion process. Diffusion-TS [293] overcomes this by applying the diffusion process after decomposition, thereby preserving the characteristics of each component more effectively. Through this approach, the traditional limitations of time series decomposition are overcome, allowing for a more effective separation of trend and seasonal components.
@@
Non-linear Methods
While Moving Average Kernel and Downsampling are linear methods, Empirical Mode Decomposition (EMD) [105] is a non-linear and non-stationary signal analysis technique that decomposes time series data into a number of Intrinsic Mode Functions (IMFs). EMD involves finding local extrema in the signal and using linear interpolation for the upper and lower envelopes, and by subtracting their average, it extracts high-frequency components as IMFs iteratively. The final residual signal, also known as the residue, reveals the long-term trend. EMD is advantageous for processing non-linear and non-stationary time series data due to its adaptive nature, effectively separating and analyzing complex patterns and changes in the time series. EMD-BI-LSTM [172] proposes a method that combines EMD with a conventional Bi-LSTM model for forecasting electric load time series data. By applying EMD, the model effectively captures complex non-linear relationships, thereby enhancing forecasting performance.
However, EMD has a mode mixing problem, where signals of different frequencies are mixed within the same IMF [274]. In the signal decomposition process of EMD, IMF is generated based on local extrema, which causes difficulties in independently separating components for signals with sudden changes, such as intermittent signals. This leads to a failure in frequency separation, which not only reduces prediction performance but also impairs interpretability. Ensemble Empirical Mode Decomposition (EEMD) [271] addresses this mode mixing problem by adding noise to the original signal and repeating the decomposition process, averaging the results. Through this process, signals with different scales are properly aligned into the correct IMFs. Through successive iterations, the influence of the noise is reduced, and the components of the original signal are emphasized, providing more physically meaningful results and enabling more accurate and stable time series decomposition. [286] proposes models applying EEMD to LSTM, linear regression, and Bayesian ridge regression for forecasting.
Variational Mode Decomposition (VMD) [52] is a method that uses variational optimization to separate signals into modes with specific frequency bands. Each mode’s frequency center and bandwidth are determined through variational optimization. As the frequency ranges do not overlap, the separation is clearer and more stable. However, compared to EMD, which requires relatively simple calculations, VMD is computationally more complex due to its use of variational optimization. [193] proposes a new model for crude oil price forecasting that applies VMD. By decomposing the time series data with VMD, transforming it into images, and extracting features with CNN, this method improves forecasting performance for oil prices. It effectively utilizes the advantages of VMD by making predictions with Bidirectional Gated Recurrent Unit (BGRU).
Various decomposition methods are widely used in time series forecasting. However, decomposition methods also have some drawbacks. When the data is divided into separate components based on different rules, the model predicts each component independently. As a result, important interactions or relationships may not be considered, which can lead to underutilization of necessary information for prediction and degrade performance. Nevertheless, considering the characteristics of time series, such as periodicity, trends, and noise, decomposition remains a powerful technique. Moreover, since time series data is more difficult to intuitively interpret compared to other data types, decomposition provides interpretability, which helps increase the model’s trustworthiness.